We've tried and tried.........
Mary makes 3 investments. The investment that pays 12% is twice the amount of the account that pays 9%. The amount invested at 10% is $500 more than the amount invested at 9%. If the annual interest income is $1555, how much money is invested at each rate?
We tried making the box-
P R T = I
1. 2x .12 1
2. x .09 1
3. x+500 .10 1
We're stuck. Line #1 we got $840.00
#2 $315.00
Then it doesn't add up to 1555. Can you help us figure this out?
Thanks,
Tricia

Answer by checkley77(12844)

.12*2x+.09x+.10(x+500)1,555
.24x+.09x+.10x-50=1,555
.43x=1,555-50
.43x=1,505
x=1,505/.43
x=3,500 is the amount invested @ 9%
2*3,500=7,000 invested @ 12%
3,500+500=4,000 invested @10%
Proof:
.12*7,000+.09*3,500+.10*4,000=1,555
840+315+400=1,555
1,555=1,555